# Task Matters: Knowledge Requirements Shape LLM Responses to Context–Memory Conflict

### v2

@@ 4 Findings
Figure 3: Performance of each model on different task types. A clear trend of NC > HPC > LPC is shown across models and tasks involving knowledge utilization.
Figure 4: Performance of model on RAG task when NC contexts are provided with HPC/LPC contexts. All models show a preference for plausible contexts.
The performance of each model on each task type and context type is reported in fig.˜3. A universal trend can be observed: regardless of the tasks, all models suffer when asked to provide responses that contradict their parametric knowledge. The behavior of the models on PC examples, however, differs by task, suggesting that the effect of convincing examples varies based on the task’s expectation.
In CK tasks (fig.˜3(b)), the model is explicitly instructed to ignore its own beliefs and rely solely on the given passage. Nevertheless, every model shows a clear NC>>HPC>>LPC performance ordering, indicating that the model still relies on parametric knowledge when it is not supposed to. This aligns with prior work’s finding that models favor their parametric knowledge more than the given contextual knowledge, thus leading to hallucinations Jin et al. (2024a). This issue, if left untreated, could not only affect the overall performance but also the correctness of model-based evaluation results, which we illustrate in §4.3.
Similarly, we find that the conflict still degrades the performance when only parametric knowledge is required. fig.˜3(c) examines model performance under settings where only parametric knowledge is needed. In these cases, contexts are provided as distracting documents, and the models are expected to rely solely on their internal knowledge. We observe a consistent degradation in accuracy when the input includes conflicting contextual passages (either HPC or LPC) compared to NC instances. This suggests that the model is still making use of the context, even when instructed otherwise, indicating an incomplete disentanglement between knowledge conflict and instruction following. Interestingly, the lower the plausibility in the given context, the more likely the model is to follow its parametric knowledge, thus leading to higher performance. This suggests that, although plausible contexts can lead to more context reliance, they can also be harmful when the underlying task requires less context reliance. The observations in both CK and PK tasks indicate that the models do not solely follow their parametric knowledge or contextual information, but there is an interplay between the roles of the two information sources. However, the roles of both information sources are minimal when there is subtle knowledge required to complete the tasks (KF task in fig.˜3(a)).
Models favor the more plausible passage when two passages compete.
Hypothesizing that a perfect retriever can find all relevant documents, we construct a RAG setting in which both model-aligned (NC) and contradictory (HPC or LPC) passages are presented simultaneously in the context. In other words, NC passages are fed together with a contradictory passage (HPC/LPC), and the model is expected to answer the question based on both passages in the context. The result is shown in fig.˜4. Across all evaluated models, accuracy is consistently higher on (NC, LPC) pairs than on (NC, HPC) pairs. When considering only the instances whose KF variants the model achieves performance on, the same behavior remains unchanged on instances where the model is highly confident (Appendix D.1), confirming our findings in this section. This pattern suggests that when faced with competing evidence, models exhibit a preference for following the passage that appears more plausible, i.e., the one more consistent with real-world knowledge. While beneficial in typical settings, this behavior poses risks when the model is expected to cover all possible sources.
@@ 4.1 Conflict Impairs Model Performance on Knowledge-Intensive Tasks
Figure 6: Averaged error distribution on RAG and PCK task. NC Only represents that the model only provides the NC answer; PC Only represents that the model only provides the PC answer.
@@ 4.2 Rationales and Reiteration
§4.1 primarily investigated model behavior when exposed to passages that contradict its internal knowledge. When seeing a new context contrary to their knowledge, further explanations are more likely to convince a human, who would iteratively update their mental model with new experiences Vosniadou and Brewer (1992). Similarly, Xie et al. (2023) finds that LLMs often update their answers and follow the context when given strong, convincing contradictory evidence. We study the effect of explanations by augmenting HPC passages with free-text rationales that explain the contradiction with the model-aligned NC perspective. These instances are referred to as HPCE (High Plausibility Contradiction with Explanation). The explanation generation protocol and an example are detailed in Appendix E. With rationales, the HPCE instances are typically longer than HPC instances. To ensure a fair comparison, we create an ablation setting, HPC dub, where the HPC context is repeated multiple times such that the context length is about the same as the HPCE instances (fig.˜5).
Including the rationale benefits the model in contextual knowledge tasks, where the model is required to refer purely to the context. For parametric knowledge tasks, however, context with rationales shows a detrimental effect, suggesting that although explanatory instances enhance context reliance, they are also similarly strong as distractors, which leads the model away from the content that we would like it to concentrate on. Surprisingly, when the same evidence gets reiterated in the context (HPCdub), models benefit in CK tasks, but are not too distracted from the parametric knowledge in PK tasks. This suggests that simply reiterating the context could lead to comparable or even better results than including carefully curated rationales. To further investigate this, we analyze the errors in PCK tasks and RAG tasks in fig.˜6 and Appendix D. We find that, although simply reiterating context leads to fewer errors, among all the mistakes the model makes, it is more likely to provide only the answers that appear more frequently, resonating with prior works that find models exhibit majority bias Gupta et al. (2023). Such behavior presents both advantages and limitations. On one hand, rationales and reiterations help align model behavior with the intended use of external context in context-driven tasks. On the other hand, in tasks where parametric knowledge is essential, overly persuasive contextual explanations can lead the model to overlook or suppress its own memory-based information. This suggests that the reliance on context should be task-dependent, highlighting the potential of manual context-memory balancing methods by controlling attention heads Jin et al. (2024b).
@@ 4.3 Conflict leads to unreliable judge LLMs
LLMs have been increasingly used as evaluators in settings where generated responses must be judged along several criteria, including whether the response contains the same information as a ground truth answer Zheng et al. (2023); Liu et al. (2023); Ru et al. (2024); Chen et al. (2025). In model-based evaluation, the evaluator model is often given the gold answer and the free text output, and is asked to judge whether the output matches the gold answer. One of the source data of our dataset, WikiContradict Hou et al. (2024), employs a language model as a judge to decide whether the free-text answer aligns with the gold answer. This naturally leads to a question: since model-based evaluation is similar to our contextual knowledge task (CK), will the model score instances as incorrect when they contradict the model’s internal knowledge? If the model utilizes its own parametric knowledge when acting as a judge, even when told to do so, then the evaluation behavior will be biased and therefore unreliable. To answer this question, we create a free generation version of our diagnostic framework following Hou et al. (2024) and perform a small-scale human annotation on 50 examples. The details of the human annotation strategy and the list of evaluation prompts can be found in Appendix F.1. We find that the averaged Cohen’s $\kappa$ Landis and Koch (1977) between the evaluator model (GPT-4o) and human annotator is 0.79 (substantial agreement), which is significantly lower than $\kappa=0.90$ (almost perfect agreement) between the human annotators. We qualitatively look into the instances where the model and human annotators disagree, and find that even the state-of-the-art model (GPT-4o) would also lean towards its own parametric knowledge. An example of such an instance is presented in fig.˜7, where GPT-4o fails to adhere to the instruction and refuses to grade an output that is contextually correct but factually incorrect as correct. One may consider employing a conflict alleviation technique to enforce stronger context reliance, but blindly following the context could also increase the risk of prompt injection Perez and Ribeiro (2022); Greshake et al. (2023). Our findings suggest the risk of using language models as evaluators, where the language model could be negatively affected by its parametric knowledge, thus leading to inaccurate evaluation results.
@@ 5 Conclusion
LLMs must constantly arbitrate between what is written in the prompt and what is stored in their parameters. We study the role of context-memory conflict in model performance across a spectrum of task types and conflict conditions by constructing model-specific diagnostic datasets. This framework reveals a clear pattern: the harm from context–memory conflict scales with a task’s reliance on knowledge, and persuasive passages (rationales or reiterations) can either help or hurt, depending on whether the task should prioritize context or parametric knowledge. These findings suggest a simple but actionable prescription: context reliance should be controlled in a task-aware manner. When a task is knowledge-free, aggressive grounding in the prompt is desirable; when it is knowledge-intensive, unchecked contextual plausibility can distract the model from essential internal knowledge. Concretely, future systems could manually modulate the model’s attention pathways Jin et al. (2024b) to balance the two sources of knowledge based on the task requirements, while the exact requirement of knowledge involvement by task still require future study. Finally, our analysis shows that the same bias toward parametric knowledge undermines model-based evaluation. Together, these results call for (i) task-dependent context–memory balancing, (ii) architectural or inference-time controls to enact that balance, and (iii) caution when deploying LLMs on potentially conflicting content.
@@
The creation of our diagnostic instances relies on LLMs, which may introduce biases, hallucinations, or artifacts that do not reflect real-world task distributions. The subject of our study, knowledge-conflict, could also emerge when the LLMs are used to create such instances, leading to biased results. Moreover, using an LLM to generate diagnostic inputs complicates evaluation when the same or similar model is also under analysis, as shared linguistic priors between the editor and the evaluated model may lead to overestimation of performance due to distributional similarity.
Disentangling memory and instruction following.
In many NLP studies, knowledge is usually framed as factual or propositional content Lewis et al. (2020); Chen et al. (2022); Meng et al. (2022a); Mallen et al. (2023). We loosely define extractive QA as a knowledge-free task. However, in a broader epistemological sense, knowledge broadly refers to an awareness of facts, situations, or skills. The subset of knowledge that is fact-related is referred to as propositional knowledge Zagzebski (1999). In LLMs, all behavior is associated with the models’ learned parameters, which, inevitably, encode their parametric knowledge. Prior work attempted to locate and modify specific factual beliefs embedded within a model’s parameters Meng et al. (2022a, b); Armengol-Estapé et al. (2024). However, modifying propositional knowledge can also lead to unintended alterations in the model’s behavior Meng et al. (2022a). Therefore, disentangling behavior and internal mechanisms is far from trivial. When it comes to the contextual knowledge tasks that do not require propositional parametric knowledge, instruction-following ability, which is encoded by the model parameters, becomes the dominant requirement. Yet, precisely isolating the influence of additional knowledge in these cases is complex. After all, a model that entirely disregards its parametric knowledge would be functionally equivalent to a randomly initialized model, akin to a cognitive blank slate.
Acknowledgement
The authors thank Niyati Bafna, Krithika Ramesh, and members of the Dredze Lab for their helpful feedback.
@@ Appendix A Parametric Knowledge Query
We query for the parametric knowledge with multiple prompts. For a single instance $(q_{i},\{a_{i1},a_{i2}\},\{c_{i1},c_{i2}\})$ in dataset $D_{\text{orig}}=\{(q_{i},\{a_{i1},a_{i2}\},\{c_{i1},c_{i2}\}),i\in[1,N]\}$, we prompt the model to confirm whether they believe the answer to $q_{i}$ is $a_{i1}$ or $a_{i2}$. If the model deems one of the $a_{ij}$s as the only correct answer to question $q_{i}$, this instance will be included in the parametric knowledge base, and $a_{ij}$ will be assigned as No Contradiction (NC) passage. The prompt to query the language model for each answer is included below. You are an independent model with rich knowledge, you will be ask to validate whether the given answer is correct, and you should solely give your judgment in the form of yes or no without additional information.
@@
Appendix B Prompts
B.1 Evidence Creation Prompts
We generate LPC and HPCE examples with GPT-4o, after a few round of prompt engineering. The final prompts used for evidence creation are shown in Figure 8.
The resulting evidence is then passed to plausibility examination. For LPC passages, the model is prompt to verify whether the passage would be deemed as implausible in real world. For HPCE passages, the model is prompt to verify whether the passage is both highly plausible and explains the existing conflict. The final prompt is included in Figure 9.
[…]
Figure 8: Final prompt for evidence creation.
B.2 Task-Annotation Prompts
[…]
As the base dataset we start with already provided answer to the questions, we only need to annotate the task under the case of knowledge free setting. We pose the knowledge free tasks as extractive question-answering task, requiring the model only to copy over the answer (Figure 12). Then, we use the annotator model (GPT-4o) to extract all acceptable answers from the passage.
B.3 Validation Prompts
The final data will be passed to language model for validation (validation in Figure 2). The final prompts used for validation is included in Figure B.3.
Appendix C Task Examples

### v3

@@ 4 Findings
Figure 3: Performance of each model on different task types. A clear trend of NC > HPC / LPC is shown across tasks involving knowledge utilization.
@@
Figure 5: Averaged error distribution on RAG and PCK tasks. NC Only represents that the model only provides the NC answer; PC Only represents that the model only provides the PC answer; Both Wrong represents the case where the model provides neither PC or NC answer.
The performance of each model on each task type and context type is reported in Figure˜3. A universal trend can be observed: in all but Knowledge-Free tasks, all models suffer when asked to provide responses that contradict their parametric knowledge.
In CK tasks (Figure˜3(b)), the model is explicitly instructed to ignore its own beliefs and rely solely on the given passage. Nevertheless, every model shows a clear NC>>HPC//LPC performance ordering, indicating that the model still relies on parametric knowledge when it is not supposed to. This aligns with prior work’s finding that models favor their parametric knowledge more than the given contextual knowledge, thus leading to hallucinations Jin et al. (2024a). This issue, if left untreated, could not only affect the overall performance but also the correctness of model-based evaluation results, which we illustrate in §4.3.
Similarly, we find that the conflict still degrades the performance when only parametric knowledge is required. Figure˜3(c) examines model performance under settings where only parametric knowledge is needed. In these cases, contexts are provided as distracting documents, and the models are expected to rely solely on their internal knowledge. We observe a consistent degradation in accuracy when the input includes conflicting contextual passages (either HPC or LPC) compared to NC instances. This suggests that the model is still making use of the context, even when instructed otherwise. To verify that this failure is not due to limitations in the model’s instruction-following ability, we vary the strength of instructions to control the absolute performance on each setting, and the results suggest that while the absolute performance may vary, the relative trends between each context type remain unchanged (Appendix F). Unsurprisingly, the lower the plausibility in the given context, the more likely the model is to follow its parametric knowledge, thus leading to higher performance. This suggests that, although plausible contexts can lead to more context reliance, they can also be harmful when the underlying task itself expects less context reliance. Taken together, the results from both CK and PK tasks show that the impact of context–memory conflict highly depends on the task’s intended knowledge reliance: the same contextual information can be either helpful or harmful depending on whether the task is designed to prioritize contextual grounding or parametric knowledge. However, the roles of the conflict are minimal when there is subtle knowledge required to complete the tasks (KF task in Figure˜3(a)), and some models even perform slightly better on LPC examples. Because KF primarily involves copying from the passage, LPC instances may reduce ambiguity by providing context that is clearly differentiated from the world’s factual knowledge.
Model HPC LPC HPC – LPC
OLMo2-7B 43.3 46.3 -3.0
OLMo2-13B 60.9 65.8 -4.9
MistralMoE-7B 41.2 54.4 -13.2
Qwen2.5-7B 60.2 67.1 -6.9
Qwen2.5-14B 59.1 61.0 -1.8
Table 1: Performance of the model on the RAG task.
More plausible $\neq$ Higher reliance
Hypothesizing that a perfect retriever can find all relevant documents, we construct a RAG setting in which both model-aligned (NC) and contradictory (HPC or LPC) passages are presented simultaneously in the context. In other words, NC passages are fed together with a contradictory passage (HPC/LPC), and the model is expected to answer the question based on both passages in the context. The result is shown in Table˜1. Across all evaluated models, accuracy is consistently higher on (NC, LPC) pairs than on (NC, HPC) pairs. Similarly, in CK and PCK tasks, models do not show a strong preference for higher plausibility passages. In contrast, for PK tasks, lower-plausibility passages enable models to rely more on their parametric knowledge rather than follow the context. This behavior is counterintuitive. Prior work suggests that language models tend to trust text that appears more credible or plausible Chen et al. (2024), whereas our results indicate that such a heuristic does not strictly hold in the presence of explicit knowledge conflict. Moreover, when considering only the instances whose KF variants the model achieves performance on, the same behavior remains unchanged on instances where the model is highly confident (Appendix 5), confirming our findings in this section.
@@ 4.2 Rationales and Reiteration
§4.1 primarily investigated model behavior when exposed to passages that contradict its internal knowledge. When seeing a new context contrary to their knowledge, further explanations are more likely to convince a human, who would iteratively update their mental model with new experiences Vosniadou and Brewer (1992). Xie et al. (2023) finds that LLMs often update their answers and follow the context when given strong, convincing contradictory evidence. We study the effect of explanatory evidence by augmenting HPC passages with free-text rationales that explain the contradiction with the model-aligned NC perspective. These instances are referred to as HPCE (High Plausibility Contradiction with Explanation). The explanation generation protocol and an example are detailed in Appendix G. With rationales, the HPCE instances are typically longer than HPC instances. To ensure a fair comparison, we create an ablation setting, HPC dub, where the HPC context is repeated multiple times such that the context length is about the same as the HPCE instances (Figure˜4).
@@
Including the rationale benefits the model in CK tasks, where correct predictions require strict reliance on the context. In contrast, for PK tasks, rationales have a detrimental effect: while explanatory instances increase context reliance, they can also act as strong distractors that pull the model away from its parametric knowledge. Interestingly, when the same evidence is reiterated in the context (HPCdub), models improved CK tasks without being overly distracted from the parametric knowledge in PK tasks. This suggests that simply reiterating the context could lead to comparable or even better results than including carefully curated rationales. To better understand this effect, we analyze the errors in PCK tasks and RAG tasks in Figure˜5 and Appendix E. We find that, although reiteration reduces overall errors, the remaining mistakes disproportionately favor answers that appear more frequently in the context. While this behavior partly reflects a form of majority bias, where models preferentially produce answers that appear more frequently in the context Gupta et al. (2023), it also suggests a deeper connection to prior findings from the language model security literature. Specifically, studies on prompt-based attacks show that reiteration alone can substantially amplify a model’s internal belief or response preference. This explains why reiterated context can be as influential as explicit rationales and is beneficial in context-driven tasks. However, in tasks that require parametric knowledge, overly persuasive contextual signals can suppress the memory and lead the model away from the correct solution. Together, these observations underscore that context reliance is inherently task-dependent, and motivate explicit context–memory balancing mechanisms, such as selectively controlling attention to contextual tokens Jin et al. (2024b); Li et al. (2025) or proxy-model steering Wang et al. (2025b).
@@ 4.3 Conflict leads to unreliable judge LLMs
LLMs have been increasingly used as evaluators in settings where generated responses must be judged along several criteria, including whether the response contains the same information as a ground truth answer Zheng et al. (2023); Liu et al. (2023); Ru et al. (2024); Chen et al. (2025). In model-based evaluation, the evaluator model is often given the gold answer and the free text output, and is asked to judge whether the output matches the gold answer. One of the source data of our dataset, WikiContradict Hou et al. (2024), employs a language model as a judge to decide whether the free-text answer aligns with the gold answer. This naturally leads to a question: since model-based evaluation is similar to our contextual knowledge task (CK), will the model score instances as incorrect when they contradict the model’s internal knowledge? To answer this question, we create a free generation version of our diagnostic framework following Hou et al. (2024) and perform a small-scale human annotation on 50 examples. The details of the human annotation strategy and the list of evaluation prompts can be found in Appendix H.1. We find that the averaged Cohen’s $\kappa$ Landis and Koch (1977) between the evaluator model (GPT-4o) and human annotator is 0.79 (substantial agreement), which is significantly lower than $\kappa=0.90$ (almost perfect agreement) between the human annotators. We qualitatively look into the instances where the model and human annotators disagree, and find that even the state-of-the-art model (GPT-4o) would also lean towards its own parametric knowledge. An example of such an instance is presented in Figure˜6, where GPT-4o fails to adhere to the instruction and refuses to grade an output that is contextually correct but factually incorrect as correct. One may consider employing a conflict alleviation technique to enforce stronger context reliance, but blindly following the context could also increase the risk of prompt injection Perez and Ribeiro (2022); Greshake et al. (2023). Our findings suggest the risk of using language models as evaluators, where the language model could be biased by its parametric knowledge, thus leading to inaccurate evaluation results.
@@ 5 Conclusion
LLMs must balance parametric and contextual knowledge, yet how they do so under conflict depends on the task at hand. In this work, we show that the impact of context–memory conflict is task-dependent: conflicts have little effect on tasks requiring minimal knowledge utilization, but substantially impair performance on knowledge-intensive tasks. By holding underlying knowledge constant while varying task formulations, our framework offers a unified explanation for previously fragmented findings in the literature.
Our results further highlight that increasing context reliance through strategies such as reiteration could be beneficial, but can also be harmful when parametric knowledge should dominate. Moreover, we show that task-dependent knowledge conflict undermines the reliability of model-based evaluation, as LLM judges may be systematically biased by their own parametric knowledge. Together, these findings suggest that both blindly enforcing context adherence and trusting only parametric knowledge can be detrimental, motivating mechanisms that dynamically balance contextual and parametric knowledge at inference time (Wang et al., 2025b; Li et al., 2025).
@@
Our diagnostic instances are partly generated or edited with the assistance of LLMs, which may introduce biases, hallucinations, or artifacts that do not reflect real-world task distributions. The subject of our study, knowledge conflict, could also emerge when LLMs are used to edit such instances, even when the original context was authored by humans. Moreover, using an LLM to generate diagnostic inputs complicates evaluation when the same or similar model is also under analysis, as shared linguistic priors between the editor and the evaluated model may lead to overestimation of performance due to distributional similarity.
Definition of Knowledge.
In NLP studies, knowledge is usually framed as factual or propositional content Lewis et al. (2020); Chen et al. (2022); Meng et al. (2022); Mallen et al. (2023). We loosely define extractive QA as a knowledge-free task. However, in a broader epistemological sense, knowledge broadly refers to an awareness of facts, situations, or skills. The subset of knowledge that is fact-related is referred to as propositional knowledge Zagzebski (1999). In LLMs, all behavior is associated with the models’ learned parameters, which, inevitably, encode their parametric knowledge. This unsettled and multifaceted definition of knowledge poses challenges for reliably quantifying the degree of knowledge involvement in tasks, complicating efforts to systematically study knowledge conflicts.
Ethical Considerations
This work analyzes how large language models resolve conflicts between contextual information and parametric knowledge. While our framework is diagnostic and does not introduce new model capabilities, it highlights several ethical considerations related to the evaluation and deployment of models.
Our findings show that LLMs used as evaluators can be systematically biased by their own parametric knowledge when judging outputs that conflict with the provided context, raising concerns about the reliability of model-based evaluation. In addition, strategies that indiscriminately increase context reliance may be harmful for tasks that require parametric knowledge and may increase vulnerability to prompt injection.
The datasets constructed in this work are derived from existing human-cleaned benchmarks and synthetic edits, and do not involve personal, sensitive, or human-subject data. Overall, our results emphasize the need for task-aware evaluation and deployment practices that explicitly account for differing knowledge requirements across tasks.
@@ Appendix A Parametric Knowledge Query
We query for the parametric knowledge with multiple prompts. For a single instance $(q_{i},\{a_{i1},a_{i2}\},\{c_{i1},c_{i2}\})$ in dataset $D_{\text{orig}}=\{(q_{i},\{a_{i1},a_{i2}\},\{c_{i1},c_{i2}\}),i\in[1,N]\}$, we prompt the model to confirm whether they believe the answer to $q_{i}$ is $a_{i1}$ or $a_{i2}$. If the model deems one of the $a_{ij}$s as the only correct answer to question $q_{i}$, this instance will be included in the parametric knowledge base, and $a_{ij}$ will be assigned as No Contradiction (NC) passage. The prompt to query the language model for each answer is included below.
You are an independent model with rich knowledge. You will be asked to validate whether the given answer is correct, and you should solely give your judgment in the form of yes or no without additional information.
@@
Appendix B Verification of Data Validity
In the final stage of the diagnostic data creation flow (Validation in Figure˜2), all instances are verified by two language models and subsampled for human verification. The number of instances removed during this process is reported in Table˜3.
Model Verification
Each (question, evidence, answer) triple was validated by GPT-4o and DeepSeek-R1-Distill-Llama-70B Guo et al. (2025). Both models were required to answer the question using the provided gold evidence. An instance was retained only if both predictions matched the annotated answer.
Human Verification
To further ensure data quality, a subset of model-verified instances was manually reviewed. For each test model (OLMo, Mistral, and Qwen), 30 randomly selected instances were manually reviewed. Annotators verified (1) whether the evidence supported its answer, and (2) whether the gold answer was correct for the corresponding task (e.g., CK answer for CK task; PK answer for PK task). All sampled instances passed human verification, likely reflecting the strictness of the preceding model-based filter. The annotators were graduate students in NLP with prior annotation experience. They were compensated at standard research assistant rates.
Appendix C Dataset Statistics
Because each model encodes different parametric knowledge, the resulting diagnostic datasets differ across models. The overall statistics are reported in Table˜2. Each question is paired with four evidence conditions (NC, HPC, HPCE, LPC), so the effective dataset size is four times the number of base instances.
Model # Instances
Mistral-7B 2,893
OLMo2-7B 177
OLMo2-13B 456
Qwen2.5-7B 6,217
Qwen2.5-14B 4,250
Table 2: Number of instances of the resulting data for each model.
Stage Mistral OLMo2-7B OLMo2-13B Qwen2.5-7B Qwen2.5-14B
Parametric knowledge querying (# dropped) 10,752 13,487 13,207 7,432 9,379
Model verification (# dropped questions) 22 4 4 18 38
Model verification (# dropped instances) 88 16 16 72 152
Final (# questions) 2,893 176 456 6,217 4,250
Table 3: Instances removed during each stage of data creation and filtering. Each question is associated with four evidence types, so the total number of evidence–answer pairs equals four times the final question count; the number of dropped instances scales accordingly.
Appendix D Task Examples